### v1

# Ultrafast Band Structure Dynamics in Bulk 1T-VSe

Abstract
Complex materials encompassing different phases of matter can display new photoinduced metastable states differing from those attainable under equilibrium conditions. These states can be realized when energy is injected in the material following a non-equilibrium pathway, unbalancing the unperturbed energy landscape of the material. Guided by the fact that photoemission experiments allow for detailed insights in the electronic band structure of ordered systems, here we study bulk 1T-VSe in its metallic and charge-density-wave phase by time- and angle-resolved photoelectron spectroscopy. After near-infrared optical excitation, the system shows a net increase of the density of states in the energy range of the valence bands, in the vicinity of the Fermi level, lasting for several picoseconds. We discuss possible origins as band shifts or correlation effects on the basis of a band structure analysis. Our results uncover the possibility of altering the electronic band structure of bulk 1T-VSe for low excitation fluences, contributing to the understanding of light-induced electronic states.
I introduction
Layered transition-metal dichalcogenides (TMDCs) are attracting particular attention since they exhibit a variety of physical properties such as spin- and valley-coupled two-dimensional superconductivity, exciton Hall effect and exotic charge-density-wave (CDW) transitions.[1]
Path breaking works on light induced CDWs in compunds as VO[2, 3] and LaTe [4] triggered the search for similar effects in other types of materials, as TMDCs.
Photoexcitation can drive TMDC-CDW systems in metastable states surviving for several picoseconds, as discussed for the examples of 1-TiSe[5, 6], 1-TaSe[7] and 1-VSe[8].
Origins for these states are attributed to phonon bottle-neck effects and optically induced phase transitions from the CDW phase into a metallic phase.
Duan et al. discovered that for TiSe an optically induced metastable metallic phase can be induced far below the CDW phase transition temperature, showing different properties than the equilibrium normal phase.[6]
As the origin the optical triggering of a halted motion of the atoms via coherent electron-phonon coupling processes was discussed.
In the case of 1T-VSe instead the phonon bottle-neck effects were discussed after optical excitation in its metallic phase.[8]
Recent interest has been drawn to the monolayer limit of VSe due to the enhanced CDW transition temperature and enlarged band gap.[9, 10, 11, 12]
Biswas et al.[9] report on ultrafast light induced insulator-metal transition in 2D VSe.
However, there are still open questions regarding the bulk properties of the material to be answered to gain a profound understanding of the physical mechanisms that determine the materials properties and phase transitions.
With this work we contribute to the understanding of the optically induced metastable state of 1T-VSe by performing time- and angle-resolved photoemission (tr-ARPES) experiments.
We excite the material with 1.55 eV pump pulses and probe at its point with 200 fs time and 50 meV energy resolution.
Comparing results acquired at three different sample temperatures (well above, around and well below the CDW transition temperature of T110 K[13]), we also address the influence of the initial phase of the system on the observed metastable modification.
ARPES gives access to the valence band structure of a material, which in the case of 1T-VSe is given by the vanadium 3d (V) and selenium 4p (Se) bands.[14, 15, 8]
Beyond the equilibrium band structure determination, tr-ARPES allows studying the material response to an optical stimulus that drives the system far from equilibrium.
This technique allows to study band-specific carrier dynamics which reveal e.g. information about the electron-electron and electron-phonon interactions in the system under investigation.
Moreover, Tr-ARPES has been shown to be an adequate technique to study light-induced metastable states in the class of TMDC materials and beyond.[8, 16, 6, 5, 7]
In this work we are presenting band structure measurements on bulk 1T-VSe for different probe photon polarizations, which allows us to experimentally disentangle the contributions of the vanadium and selenium bands in the valence band structure.
We furthermore use the selective access to bands with different orbital character offered by pulses of different polarization.
In Figs. 1a,b we show the layered crystal structure of bulk 1T-VSe and its hexagonal Brillouin zone (BZ).
The two photon energies used in this work, namely 6.2 eV and 10.8 eV, probe the three dimensional band structure in the AHL and KM plane, respectively.
II Methods
II.1 Photoelectron spectroscopy
ARPES measurements at equilibrium were performed using a photon energy of 10.8 eV.
The angle of incidence of the laser beam is 30 with respect to the surface normal.
The polarization of the 10.8 eV photons is changed by altering the polarization of the incoming beam in the last harmonic upconversion step, which is a third-harmonic generation stage performed in a cell filled with Xe gas. More details about the setup are reported in Ref. [17].
Due to the measurement geometry -polarized light always keeps a component parallel to the sample surface, whereas -polarized light probes only states with in plane components and cannot address states with a nodal plane parallel to the sample surface.
Tr-ARPES experiments were performed using the 800 nm fundamental beam of an amplified short pulse Ti:sapphire laser system as a pump beam and its 4th harmonic as a probe.
The polarization of the probe beam is varied by a wave plate.
All data were acquired with a SPECS PHOIBOS 225 hemispherical electron analyzer, equipped with a delay line detector by Surface Concept. We recorded all data in the wide-angle mode of the spectrometer.
Clean sample surfaces were prepared by cleaving in situ a bulk VSe crystal, purchased from HQ Graphene, using a kapton tape.
The background pressure of the ultra-high vacuum chamber, in which we were cleaving the samples amounted to mbar, while the measurements were performed at a background pressure of mbar.
The orientation of the sample on the sample holder was determined by investigating its low-energy electron diffraction pattern.
II.2 Density functional theory calculations
Density functional theory calculations were performed using the Vienna ab initio package (VASP)[18, 19, 20] with the projector augmented wave method[21, 22]. Freestanding films were modeled with a 16 Åvacuum gap for the supercell. A plane-wave energy cutoff of 320 eV and a 16x16x10 k-mesh were adopted for the films. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) functional[23] was used; the bulk lattice constants are Åand Å. Spinorbit interactions were incorporated throughout the computations.
III results
III.1 Polarization dependent ARPES at equilibrium
The valence band structure of 1T-VSe originates from the V and the Se bands.
According to band structure calculations, these bands show different dispersion along the A direction of its hexagonal BZ.[24, 8, 25, 26]
Whereas four of those bands cross at the point of the BZ, at the A point these bands are separated from each other.
By performing ARPES using different photon energies and light polarizations, these bands are distinguishable by their different dispersion in the three dimensional momentum space and orbital band character.
DFT band structure calculations reported clear changes in the orbital character of the bands, despite strong hybridizations in certain regions of the BZ.[8, 24]
In the following we will compare ARPES measurements for different light polarizations to demonstrate that different parts of the valence band structure can be emphasized with - and -polarized light.
Figure 1c shows the valence band structure of bulk VSe, as probed at room temperature with photons of an energy of 10.8 eV.
Considering the dispersion of the valence bands and the inner potential of 16 eV reported by Falke et al. [25], we conclude that for free electron final states we are probing the sample band structure close to the A-point of the BZ.
We map the photoemission intensity as a function of binding energy with respect to the Fermi level () and electron momentum () along the AL and AH direction, respectively.
The uppermost panel shows the result acquired with -polarized light.
We distinguish two bands, one positioned at the Fermi level with a flat dispersion along AH and a significantly steeper gradient in AL direction (A) and a second band, which shows a steep dispersion and evolves rather similar in both high symmetry directions of the BZ (B).
Next we observe a shoulder indicating a third band with a larger effective mass than B.
Our experimental results for -polarized light are in qualitative agreement with previous ARPES studies at a photon energy of 11 eV presented by Falke et al.[25]
In the middle panel we repeated the same measurement with -polarized light and resolved the third band (C), that was barely visible with -polarized light.
Mapping the intensity difference between the two measurements, in the bottom panel we distinguish the bands with a significant out-of plane component (A and B) from the one with pure in-plane components (C).
Comparing our experimental results with our DFT band structure calculations, we can identify all three bands that are predicted to show up in the investigated part of the BZ.
Band positions as well as dispersion are remarkably well described.
In agreement with calculations reported by Strocov et al. [15] and Majchrzak et al. [8] we attribute band A to the V band.
The orbital character of band B is controversial.
According to Strocov et al. it corresponds to the Se state, whereas the band structure calculations reported by Majchrzak et al. show along AH two bands rather close to each other and strongly hybridizing leading to a more -like character close to the Fermi level and changing with increasing momentum.
The dispersion and symmetry of band C along K agrees with the calculation reported by Strocov et al. for the Se.
In our data we see a clear dependence on the polarization of the incident light, which leads to the conclusion that bands B and C have different band character, which supports our interpretation that band B is derived from the Se electrons and band C from the Se ones.
Changing the probe photon energy to 6.2 eV, we move from the A plane to the plane of the BZ, if we consider the inner potential V of 16 eV as determined by Falke et al. [25] from low photon energy synchrotron measurements in the range from 9 to 17 eV. This inner potential is much larger than the value of 7.5 eV derived from synchrotron measurements at energies in the ranges of 65 to 160 eV (cf. Ref. [8], supporting material) and 845 to 960 eV (cf. Ref. [15]). The discrepancy between the two effective inner potential values for different incident photon energies could base on final state effects. Smaller variations of V can be observed in narrower photon energy ranges.[27] However, we assume that our photon energy of 6.2 eV is close enough to the photon energies used by Falke et al. and conclude that we probe very close to the point of the BZ.
Note that this consideration does not include final state effects due to resonant excitation with our probe energy. The 6.2 eV photon energy reduces significantly our field of view in momentum space and binding energies. As we show in Fig. 2, we observe photoemission intensity in a binding energy range of eV. From these data we can deduce the sample work-function from the low energy cut-off () of the energy distribution curves (EDCs). We find eV, which is close to the value of eV reported by Claessen et al.[28] and confirms a clean and good quality surface. The photoemission intensity increases towards the Fermi-level. We observe a slight dispersion of the intensity maximum along the momentum axis . Overlaying our band structure calculations, we find that three of the four valence bands are positioned in the binding energy range where we observe enhanced photoemission intensity. Furthermore these three bands show comparable dispersion in our momentum space field of view. The fourth band instead is clearly separated from the region where we measure highest intensity. We mark the positions of the bands with black solid lines in the EDC shown at the right side of Fig. 2a, which is taken at the point (cf. integration window indicated by the dashed grey lines on top of the intensity distribution map). With the given energy resolution we cannot distinguish the different bands crossing in this energy and momentum region. Considering the polarization difference map we discussed above for 10.8 eV measurements, we obtain light blue color where the fourth band is predicted and red color in the range of the two uppermost bands that are strongly hybridizing (cf. Fig. 2b). The band with the highest binding energy shows Se character at the -point according to the calculations shown by Majchrzak et al.[8], whereas the other three are strongly hybridizing, so that the two -bands are getting significant -character. This is in contrast to the calculations published by Jolie et al.[24], but fits to our observation comparing different probe polarizations.
III.2 Pump-probe spectroscopy
For investigating optically induced changes in the electron density of states in the vicinity to the Fermi level, we performed pump-probe experiments using 1.55 eV pump photon pulses with a fluence of mJ/cm.
According to our band structure calculations shown in Fig. 3a, there are selective values along M, AL and AH directions where resonant vertical transitions from bands with predominantly -like character (grey) to -like bands (black, blue) are predicted to be possible.
The unoccupied bands around 1 eV above the Fermi level are assigned to -like character in agreement with Ref. [24].
When pumping, we initially observe a decrease in the photoemission intensity below the Fermi level, indicating a depopulation of the bands (cf. the pump/non-pumped difference map shown in Fig. 3b) and an increase of the photoemission intensity just above the Fermi level.
This behaviour is expected for a broadening of the Fermi edge due to an increase of the electronic temperature by the pump pulse.
Indeed we are observing an almost symmetric shape of the difference EDC integrated in a small window around the -point at the delay of temporal overlap between pump and probe pulses (t, cf. Fig. 3b).
At later delays, however, the difference EDC gets highly asymmetric, clearly showing a residual enhanced intensity above the Fermi level and even a net increase of the signal below the Fermi level in comparison to equilibrium conditions (cf. Fig. 3b for the EDC difference curves at and 1 ps).
Tracing the evolution of the intensity distribution as a function of pump-probe delay at and at Å (see Fig. 3c), we observe an overall increase of the photoemission intensity in the range from -0.5 eV up to the Fermi level in the whole momentum range we are investigating, surviving for several picoseconds (cf. Fig. 3c).
At the right side of the intensity difference maps in Fig. 3c we show the difference EDCs averaged over all pump-probe delays larger than 1 ps, in order to display the spectral shape of the metastable state.
At we observe an overall intensity increase, which reaches its maximum around the Fermi level.
The signal seems to drop just below the Fermi level.
We attribute this drop to a depopulation of the vanadium and selenium bands, whose central positions are marked by the magenta lines.
This assignment is getting more clear when evaluating the difference EDCs at larger values.
The drop in intensity shifts here towards higher binding energies, according to the dispersion of the bands (cf. right panel of Fig. 3c).
In Fig. 3d we show the temporal evolution of the photoemission intensity for four distinct energy positions, indicated by the colored stripes in panel c.
The grey window is chosen in the uppermost part of our measurement window, well above the Fermi level, for extracting a transient that should be influenced mainly by our pump-probe cross correlation and hence allows us to deduce t for our dataset.
The red window describes the intensity evolution above the Fermi level, whereas the blue and the green windows are chosen to correspond to the positions of the valence bands.
In comparison to the black curve, all other curves show a delayed reach of their maximum intensity change.
A delayed depopulation of the valence bands is expected due to cascading effects, since the -values for resonant excitation with 1.55 eV are just out of our field of view.
At Å the intensity change is suppressed to close to zero in the range of the upper three valence bands.
Majchrzak et al.[8] were reporting a depletion of the Se band reaching a plateau within the first 500 fs and not repopulating up to 5 ps after photoexcitation.
They were referring this to hot carriers affecting the low-energy soft phonon spectrum of the material and the electron-phonon coupling strength.
We conclude that the spectral shape of the difference EDC’s results from an intensity increase in the whole range from -0.5 eV up to the Fermi level, which is suppressed at some energies due to a long-lived depletion of the valence band population after photoexcitation.
In contrast to Majchrzak et al.,[8] we do not observe a net reduction of the photoemission intensity with respect to the equilibrium intensity due to a strong depopulation of the V band at the point that persists for several ps.
We refer this to the lower excitation fluence we are using in our experiments, which allows us to observe the net increase of the photoemission intensity.
We will discuss possible origins of this phenomenon at the end of this section.
In Fig. 4 we compare the results at room temperature (well above T) discussed so far, with measurements performed at sample temperatures of 125 K (slightly above T) and 55 K (well below T).
Here we integrated over the whole momentum range from -0.1 to 0.1 Å.
Besides the sharpness of the Fermi edge we do not observe any significant difference between the measurements acquired at sample temperatures of 300 K and 125 K.
Cooling well below T (55 K) we find a more pronounced delayed band depletion at binding energies in the order of 0.5 eV.
After 1 ps, also in this case we find the same broad-range intensity increase as above T.
Comparing the difference EDCs for the three different temperatures (cf. Fig. 4b), we observe a clear downshift of the onset of the signal suppression due to valence band depletion with decreasing sample temperature.
We interpret this observation, as well as the retarded depletion at the higher binding energies, as indications of slight changes in the valence bands dispersion when lowering the sample temperature.
We consider these changes in the band structure as related to the CDW phase and conclude that with the given pump fluence the CDW is at least partially preserved.
The optically induced intensity increase however does not seem to depend on the crystals phase.
As shown in Fig. 4c, we evaluate the total electron yield (TEY) integrated in the complete field of view we are probing with 6.2 eV photons.
We observe an increase by 1 % of the equilibrium TEY upon 800 nm pumping.
This observation is identical for starting in the normal and the CDW phase of the system and independent of the probe light polarization.
Summing up our observations, we firmly exclude the observed intensity increase being related to a closure of the CDW gap due to optical excitation, since the signal shows no temperature dependence and our measurements at 300 K are far above the critical temperature of 110 K.
We also exclude a shift of the entire band structure as being its origin, since 1.55 eV pump and HHG probe measurements published by Majchrzak et al. showed no indication for such a shift at higher excitation fluences.[8] Indeed, optically-induced band shifts should scale with the fluence.
Furthermore, we demonstrated that our positive signal decreases following the equilibrium dispersion of the selenium bands (cf. Fig. 3c).
At larger values the data of Majchrzak et al. show an indication of a drastic shift of the V band only, whose origin however is not discussed.
Since our static data (cf. Fig. 1 and Fig. 2) clearly show that the V is probed with -polarized light only, we also exclude a shift in the vanadium band position to be the origin of our signal, since we can observe it likewise with both probe polarizations.
A recent article claimed that a flat band appears in large parts of the bulk VSe BZ and can be shifted downwards via electron doping through donor adatoms adsorbed on the sample surface.[29]
They speculated that polaron formation could be a possible origin of the discussed observation.
An effect as electron doping via our pump pulses could be an explanation for our observation, however so far theoretical confirmation of flat bands or polarons in bulk VSe is lacking.
Picosecond-scale modifications of the electronic density of states after optical excitation could also result from trapping electrons at defects in the crystal structure, as it was discussed e.g. for MoS monolayers.[30]
However, in our case of incresed intensity below the Fermi level, this would mean that the defect levels are not completely occupied under equilibrium conditions.
An additional explanation of the metastable signal can be related to the light-induced suppression of electronic correlation in 1T-VSe, resulting in the recovery of additional density of states in the energy region of the vanadium 3d bands. Indeed, Biswas et al.[9] reported that significant electronic correlation is present in VSe. In particular, the fingerprint of this effect was found in the large broadening of the vanadium 3d states below the Fermi level, which could only be captured by dynamical mean field theory calculations with a large Hubbard interaction strength of the order eV. Similarly to what reported in [31], when energy is provided to the system via photoexcitation, the strength of electronic correlations can be markedly reduced, hence new electronic states in the energy region of the vanadium 3d bands, that were suppressed by localization, can emerge, similarly to the collapse of a correlation gap in photoexcited Mott insulators.
Whichever is the origin of the photoinduced metastable electronic state with enhanced density of states that we revealed, the large increase of the carrier density at the Fermi level implies a transient enhanced metallicity of the material.
IV summary and conclusion
In this study we used polarization-dependent ARPES to disentangle the contributions of different vanadium- and selenium-derived bands to the valence band structure close to the and A points of the 1T-VSe hexagonal BZ. Our results show good agreement with theoretical band structure calculations, allowing for one-to-one assignments of the probed bands. Furthermore, we used our extracted knowledge about the position and dispersion of different bands for interpreting the changes in the electronic density of states after optical pumping. We detect the formation of a metastable configuration of the density of states reached after optical pumping, leading to additional electron density of states spread over 0.5 eV below the Fermi-level, which has not been observed so far and that asks for further investigation. However, the possibility to alter the electronic bands by ultrashort light pulses in a way not attainable under equilibrium paves the way to the design of new metastable states of matter with tailored electronic properties and novel functionalities, connected to the light-induced enhanced metallicity of the system, under the form of an enhanced density of states extending from the valence-band region up to the Fermi level.
Acknowledgements.
W.B. acknowledges financial support from the German academic exchange service (DAAD) for financing a research stay at Academia Sinica in Taiwan for establishing the collaboration.References
- Manzeli et al. [2017] S. Manzeli, D. Ovchinnikov, D. Pasquier, O. V. Yazyev, and A. Kis, Nature Reviews Materials 2, 17033 (2017).
- Morrison et al. [2014] V. R. Morrison, R. P. Chatelain, K. L. Tiwari, A. Hendaoui, A. Bruhács, M. Chaker, and B. J. Siwick, Science 346, 445 (2014).
- Muraoka et al. [2018] Y. Muraoka, H. Nagao, Y. Yao, T. Wakita, K. Terashima, T. Yokoya, H. Kumigashira, and M. Oshima, Scientific Reports 8, 17906 (2018).
- Kogar et al. [2020] A. Kogar, A. Zong, P. E. Dolgirev, X. Shen, J. Straquadine, Y.-Q. Bie, X. Wang, T. Rohwer, I.-C. Tung, Y. Yang, R. Li, J. Yang, S. Weathersby, S. Park, M. E. Kozina, E. J. Sie, H. Wen, P. Jarillo-Herrero, I. R. Fisher, X. Wang, and N. Gedik, Nature Physics 16, 159 (2020).
- Hedayat et al. [2019] H. Hedayat, C. J. Sayers, D. Bugini, C. Dallera, D. Wolverson, T. Batten, S. Karbassi, S. Friedemann, G. Cerullo, J. van Wezel, S. R. Clark, E. Carpene, and E. Da Como, Phys. Rev. Res. 1, 023029 (2019).
- Duan et al. [2023] S. Duan, W. Xia, C. Huang, S. Wang, L. Gu, H. Liu, D. Xiang, D. Qian, Y. Guo, and W. Zhang, Phys. Rev. Lett. 130, 226501 (2023).
- Shi et al. [2019] X. Shi, W. You, Y. Zhang, Z. Tao, P. M. Oppeneer, X. Wu, R. Thomale, K. Rossnagel, M. Bauer, H. Kapteyn, and M. Murnane, Science Advances 5, eaav4449 (2019).
- Majchrzak et al. [2021] P. Majchrzak, S. Pakdel, D. Biswas, A. J. H. Jones, K. Volckaert, I. Marković, F. Andreatta, R. Sankar, C. Jozwiak, E. Rotenberg, A. Bostwick, C. E. Sanders, Y. Zhang, G. Karras, R. T. Chapman, A. Wyatt, E. Springate, J. A. Miwa, P. Hofmann, P. D. C. King, N. Lanatà, Y. J. Chang, and S. Ulstrup, Phys. Rev. B 103, L241108 (2021).
- Biswas et al. [2021] D. Biswas, A. J. H. Jones, P. Majchrzak, B. K. Choi, T.-H. Lee, K. Volckaert, J. Feng, I. Marković, F. Andreatta, C.-J. Kang, H. J. Kim, I. H. Lee, C. Jozwiak, E. Rotenberg, A. Bostwick, C. E. Sanders, Y. Zhang, G. Karras, R. T. Chapman, A. S. Wyatt, E. Springate, J. A. Miwa, P. Hofmann, P. D. C. King, Y. J. Chang, N. Lanatà, and S. Ulstrup, Nano Letters 21, 1968 (2021).
- Umemoto et al. [2019] Y. Umemoto, K. Sugawara, Y. Nakata, T. Takahashi, and T. Sato, Nano Research 12, 165 (2019).
- Feng et al. [2018] J. Feng, D. Biswas, A. Rajan, M. D. Watson, F. Mazzola, O. J. Clark, K. Underwood, I. Marković, M. McLaren, A. Hunter, D. M. Burn, L. B. Duffy, S. Barua, G. Balakrishnan, F. Bertran, P. Le Fèvre, T. K. Kim, G. van der Laan, T. Hesjedal, P. Wahl, and P. D. C. King, Nano Letters 18, 4493 (2018).
- Duvjir et al. [2018] G. Duvjir, B. K. Choi, I. Jang, S. Ulstrup, S. Kang, T. Thi Ly, S. Kim, Y. H. Choi, C. Jozwiak, A. Bostwick, E. Rotenberg, J.-G. Park, R. Sankar, K.-S. Kim, J. Kim, and Y. J. Chang, Nano Letters 18, 5432 (2018).
- Diego et al. [2021] J. Diego, A. H. Said, S. K. Mahatha, R. Bianco, L. Monacelli, M. Calandra, F. Mauri, K. Rossnagel, I. Errea, and S. Blanco-Canosa, Nature Communications 12, 598 (2021).
- Terashima et al. [2003] K. Terashima, T. Sato, H. Komatsu, T. Takahashi, N. Maeda, and K. Hayashi, Phys. Rev. B 68, 155108 (2003).
- Strocov et al. [2012] V. N. Strocov, M. Shi, M. Kobayashi, C. Monney, X. Wang, J. Krempasky, T. Schmitt, L. Patthey, H. Berger, and P. Blaha, Phys. Rev. Lett. 109, 086401 (2012).
- Fanfarillo et al. [2021] L. Fanfarillo, D. Kopić, A. Sterzi, G. Manzoni, A. Crepaldi, D. T. Payne, W. Bronsch, V. Tsurkan, D. Croitori, J. Deisenhofer, F. Parmigiani, M. Capone, and F. Cilento, Journal of Electron Spectroscopy and Related Phenomena 250, 147090 (2021).
- Peli et al. [2020] S. Peli, D. Puntel, D. Kopic, B. Sockol, F. Parmigiani, and F. Cilento, Journal of Electron Spectroscopy and Related Phenomena 243, 146978 (2020).
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Phys. Rev. B 48, 13115 (1993).
- Kresse and Furthmüller [1996a] G. Kresse and J. Furthmüller, Computational Materials Science 6, 15 (1996a).
- Kresse and Furthmüller [1996b] G. Kresse and J. Furthmüller, Phys. Rev. B 54, 11169 (1996b).
- Blöchl [1994] P. E. Blöchl, Phys. Rev. B 50, 17953 (1994).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, Phys. Rev. B 59, 1758 (1999).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
- Jolie et al. [2019] W. Jolie, T. Knispel, N. Ehlen, K. Nikonov, C. Busse, A. Grüneis, and T. Michely, Phys. Rev. B 99, 115417 (2019).
- Falke et al. [2021] Y. Falke, N. Ehlen, G. Marini, A. V. Fedorov, V. Y. Voroshnin, B. V. Senkovskiy, K. Nikonov, M. Hoesch, T. K. Kim, L. Petaccia, G. Di Santo, T. Szkopek, G. Profeta, and A. Grüneis, Phys. Rev. B 104, 235137 (2021).
- Henke et al. [2020] J. Henke, F. Flicker, J. Laverock, and J. van Wezel, SciPost Phys. 9, 56 (2020).
- Tamura et al. [1985] E. Tamura, R. Feder, J. Krewer, R. Kirby, E. Kisker, E. Garwin, and F. King, Solid State Communications 55, 543 (1985).
- Claessen et al. [1990] R. Claessen, I. Schafer, and M. Skibowski, Journal of Physics: Condensed Matter 2, 10045 (1990).
- Yilmaz and Vescovo [2022] T. Yilmaz and E. Vescovo, Phys. Rev. B 105, L041101 (2022).
- Wang et al. [2015] H. Wang, C. Zhang, and F. Rana, Nano Letters 15, 339 (2015).
- Cilento et al. [2018] F. Cilento, G. Manzoni, A. Sterzi, S. Peli, A. Ronchi, A. Crepaldi, F. Boschini, C. Cacho, R. Chapman, E. Springate, H. Eisaki, M. Greven, M. Berciu, A. F. Kemper, A. Damascelli, M. Capone, C. Giannetti, and F. Parmigiani, Science Advances 4, eaar1998 (2018).